OpenStudy (anonymous):

On a manufacturer's assembly line, 3% of its merchandise comes off defective. If the company has 354 defective items, how many total items came off the assembly line?

OpenStudy (whpalmer4):

3% of x = 354 \[3\%*\frac{1}{100\%} * x = 354\]\[\frac{3x}{100}=354\]Can you solve that for \(x\)?

OpenStudy (anonymous):

divide by 100 on both sides ext...?

OpenStudy (whpalmer4):

no, I would multiply both sides by 100

OpenStudy (anonymous):

oh.. then divide

OpenStudy (whpalmer4):

\[\frac{3x}{100}*100 = 354*100\]\[\frac{3x}{\cancel{100}}*\cancel{100} = 35400\]

OpenStudy (anonymous):

3*354/100=10.62?

OpenStudy (whpalmer4):

no...\[3x = 35400\]divide both sides by 3

OpenStudy (whpalmer4):

Think about it this way: 3% means 3 out of each 100. if you have 100 items, 3% is just 3. You need more than 100 times as many (you want 354), so x must be more than 100 * 100, right? 3% of 10.62 is less than 1, and the problem says there are 354 :-)